Ask your own question, for FREE!
Mathematics 12 Online
OpenStudy (anonymous):

the computer repairman is given 8 computers to fix. he knows that among them are 4 bad video cards and 3 failed hard drives. what is the probability thats the first computer he tries has a bad video card but a healthy hard drive? help ?

OpenStudy (dumbcow):

prob of bad video card = 4/8 = 1/2 prob of failed hard drive = 3/8 --> healthy hard drive = 5/8 prob of both = 1/2*5/8 = 5/16

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!